OpenStudy (anonymous):

fractions.find the sum or difference. simplify if possible. 15/20 - 9/20=

OpenStudy (anonymous):

Similar to \[\frac{15-9}{20}\]

OpenStudy (moongazer):

6/20 = 3/10

OpenStudy (moongazer):

just add the numerator and copy the denominator. cause it similar divide both sides by two to simplify.